OpenStudy (anonymous):

a rectangular lot measure is (5w+3z)cm by (5w-3z)cm. If (2w)is added to both dimensions, how is the area affected? Explain.

OpenStudy (anonymous):

using the formula (a+b)(a-b) = a² - b² original area = (5w + 3z)(5w - 3z) = (5w)² - (3z)² = 25w² - 9z²

OpenStudy (anonymous):

where is the (2w)..?..and how is the area affected..?

OpenStudy (anonymous):

on adding 2w to both we get (7w + 3z)(7w - 3z) = (7w)² - (3z)² = 49w² - 9z²

OpenStudy (anonymous):

how is the area affected?

OpenStudy (id21):

The area of 1st lot is \[(5w+3z)(5w-3z)=25w^2-9z^2\] if you add 2w for each dimension, you get:\[(5w+3z+2w)(5w-3z+2w)=(7w+3z)(7w-3z)=49w^2-9z^2\] Are will be greater by this ratio: \[\frac{49w^2-9z^2}{25w^2-z^2}\]

OpenStudy (anonymous):

new area is more by (49w² - 9z²) - (25w² - 9z²) = 49w² - 25w² - 9z² + 9z² = 24w²

OpenStudy (anonymous):

new are will be 24w² more than the original area

OpenStudy (id21):

Harkirat is right!

OpenStudy (anonymous):

@TheChariz when we r answering, hv some patience.....☺

OpenStudy (anonymous):

ok..thank you very much.. i will try to be patience..next time..